A aircraft hanger is on a airport and the hanger needs to have a fence completely around the hanger. The hanger is two times longer then it is width. There is 1000 feet of fencing material is available. What is the dimensions of the hanger and what is the dimensions will maximize the area around the hanger that will be fenced in.
Answer by jorel1380(3719)

Perimeter=2L+2W
If the length is twice the width, then;
2(2W)+2W=1000
4W+2W=1000
6W=1000
W=166 2/3
L=333 1/3
is the maximum dimensions of the fence. The area is
166 2/3 X 333 1/3=55555.556 sq. ft..